Circular No. 6439

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-1 W. Cui, Center for Space Research, Massachusetts Institute of Technol-

ogy; W. Focke, University of Maryland and Goddard Space Flight Center (GSFC); and J. Swank, GSFC, report: “The ASM lightcurve indicates that Cyg X-1 is still in the high state, and its 1.3–12-keV flux has been varying between ∼ 0.8 and 1.5 Crab since the transition (IAUC 6404). So far, twelve snapshots (public) of the source have been made with RXTE. In the PCA band (2–60 keV), the x-ray flux varied between ∼ 0.4 and 1 Crab. After an apparent ‘settling period’ (covered by the first three observations), both the energy spectrum and power-density spectrum (PDS) seem to be stabilized. The energy spectrum can be described by a model consisting of a soft-blackbody component (kT ∼ 0.3 keV) and a broken power law whose photon index breaks from ∼ 2.9 to 2.0 at ∼ 11 keV. The PDS can be characterized by a power law, with a slope of ∼ −1 (i.e., 1/f), below ∼ 16 Hz and a steeper one above, with a slope of ∼ −2. The observation made on June 17.332–17.380 UT shows the presence of quasiperiodic oscillations (QPOs) centered at ∼ 52 mHz (Lorentzian width ∼ 19 mHz), only when the source reaches a flux peak at ∼ 1 Crab. The integrated fractional rms of the QPO is ∼ 8, 11, and 10 percent in the energy bands 2–6.5, 6.5–13.1, and 13.1–60 keV, respectively. It is not clear if this QPO is related to the 40-mHz QPO reported for the low state (IAUC 5576, 5580).”

SUPERNOVA 1983ab IN ANONYMOUS GALAXY N. Samus and S. V. Antipin, Sternberg Astronomical Institute, report

the discovery by Antipin of a supernova on Moscow collection plates taken with the Crimean 0.40-m astrograph. The coordinates of the supernova measured on the discovery plates relative to GSC stars are α = 16^{h}22^{m}34^{s}.0, δ = +36^{o}45′25′′ (equinox 2000.0), which is 6′′ west and 5′′ north of the galaxy’s center. SN 1983ab is seen on 24 plates, the first being a plate exposed on 1983 Aug. 27.76 UT (at mag B = 17.07) and the last taken on Sept. 14.85 (at B = 16.86:); maximum brightness (B = 16.35 ± 0.15) occurred on Sept. 4. SN 1983ab is not present (B [17.7) on a plate taken on 1983 Aug. 16.87, nor is it visible on one taken on 1983 Sept. 27.72 (B [17.7). SN 1983T was discovered by N. Metlova on the same plates (cf. IAUC 3873).

1996 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green